Schedule of cash payments for a service company Horizon Financial Inc. was organized on February 28. Projected selling and administrative expenses for each of the first three months of operations are as follows: March $123,500 April 116,100 May 105,700 Depreciation, insurance, and property taxes represent $26,000 of the estimated monthly expenses. The annual insurance premium was paid on February 28, and property taxes for the year will be paid in June. 73% of the remainder of the expenses are expected to be paid in the month in which they are incurred, with the balance to be paid in the following month. Prepare a schedule of cash payments for selling and administrative expenses for March, April, and May. Excel Learning Systems Inc. Schedule of Cash Payments for Selling and Administrative Expenses For the Three Months Ending May 31 March April May March expenses: Paid in March $ 71,175 Paid in April $ April expenses: Paid in April Paid in May $ May expenses: Paid in May Total cash payments $ $ $
Answer #1

Answer-

HORIZON FINANCIAL INC.
SCHEDULE OF CASH PAYMENTS FOR SELLING AND ADMINISTRATIVE EXPENSES
FOR THE THREE MONTHS ENDING MAY 31
PARTICULARS MARCH APRIL MAY
$ $ $
March expenses:-
Paid in March ($123500-$26000)*73%=$71175
Paid in April ($123500-$26000)*27%=$26325
April expenses:-
Paid in April ($116100-$26000)*73%=$65773
Paid in May ($116100-$26000)*27%=$24327
May expenses:-
Paid in May ($105700-$26000)*73%=$58181
Total cash payments $ 71175 92098 82508